OpenStudy (anonymous):

the tangent to the curve x^3+x^2y+4y=1 at the point (3,-2) has slope a)-3 b)-23/9 c)-27/13 d)-11/9 e)-15/13

OpenStudy (loser66):

take derivative and plug the point in.

OpenStudy (anonymous):

The derivative can be computed in a few ways. I'll suggest two. One method is via implicit differentiation, which yields \[3x^2+2xy+x^2\frac{dy}{dx}+4\frac{dy}{dx}=0\] From here you can solve for \(\dfrac{dy}{dx}\) and plug in \(x=3\) and \(y=-2\). A slightly different way: Separate your variables to solve for \(y\), then take the derivative right away. \[\begin{align*} x^3+x^2y+4y&=1\\\\ (x^2+4)y&=1-x^3\\\\ y&=\frac{1-x^2}{x^2+4}\\\\ \frac{dy}{dx}&=\cdots \end{align*}\]